40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Jun 23 13:08:41 2018, Jon, Update, AUX, First Coherent AUX Scan of PRC Using AM Sidebands IMG_2695.jpegIMG_2698.jpegIMG_2699.jpeg
    Reply  Sat Jun 23 20:54:35 2018, Koji, Update, AUX, First Coherent AUX Scan of PRC Using AM Sidebands 
    Reply  Tue Jun 26 10:06:39 2018, keerthana, Update, AUX, First Coherent AUX Scan of PRC Using AM Sidebands TRANS.pdfREFL.pdf
Message ID: 14010     Entry time: Sat Jun 23 13:08:41 2018     Reply to this: 14011     14017
Author: Jon 
Type: Update 
Category: AUX 
Subject: First Coherent AUX Scan of PRC Using AM Sidebands 

[Jon, Keerthana, Sandrine]

Thu.-Fri. we continued with PRC scans using the AUX laser, but now the "scanned" parameter is the frequency of AM sidebands, rather than the frequency of the AUX carrier itself. The switch to AM (or PM) allows us to coherently measure the cavity transfer as a function of modulation frequency.

In order to make a sentinel measurement, I installed a broadband PDA255 at an unused pickoff behind the first AUX steering mirror on the AS table. The sentinel PD measures the AM actually imprinted on the light going into the IFO, making our measurement independent of the AOM response. This technique removes not only the (non-flat) AOM transfer function, but also any non-linearities from, e.g., overdriving the AOM. The below photo shows the new PD (center) on the AS table.

With the sentinel PD installed, we proceeded as follows.

  • Locked IFO in PRMI on carrier.
  • Locked AUX PLL to PSL.
  • Tuned the frequency of the AUX laser (via the RF offset) to bring the carrier onto resonance with the PRC.
  • Swept the AOM modulation frequency 0-60 MHz while measuring the AUX reflection and injection signals.

The below photo shows the measured transfer function [AUX Reflection / AUX Injection]. The measurement coherence is high to ~55 MHz (the AOM bandwidth is 60 MHz). We clearly resolve two FSRs, visible as Lorentzian dips at which more AUX power couples into the cavity. The SURFs have these data and will be separately posting figures for the measurements.

With the basic system working, we attempted to produce HOMs, first by partially occluding the injected AUX beam with a razor blade, then by placing a thin two-prong fork in the beam path. We also experimented with using a razor blade on the output to partially occlude the reflection beam just before the sensor. We were able to observe an apparent secondary dip indicative of an HOM a few times, as shown below, but could not repeat this deterministically. Besides not having fine control over the occlusion of the beams, there is also large few-Hz angular noise shaking the AS beam position. I suspect from moment to moment the HOM content is varying considerably due to the movement of the AS beam relative to the occluding object. I'm now thinking about more systematic ways to approach this.

 

ELOG V3.1.3-